OpenStudy (anonymous):

Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method.2x + 5y = -3 -15x - 5y = -10

OpenStudy (anonymous):

I would use elimination because the first step is already done for you

myininaya (myininaya):

First of all I would never ever choose graphing to find the solution to a set of linear equations.

OpenStudy (mertsj):

I would use elimination. Since the first equation has 5y and the second equation has -5y elimination would be convenient. I could add the two equations and eliminate the y terms.

myininaya (myininaya):

elimination sounds cool!

myininaya (myininaya):

for the reason mertsj pointed out

OpenStudy (anonymous):

I will simply substract 1 from 2... wait that is how elimination works ... isn't ? ;)

myininaya (myininaya):

just add the equations sweety

OpenStudy (anonymous):

lol yeah, subtraction is adding though :P

myininaya (myininaya):

that means you are doing this -5y-5y we are trying to do this -5y+5y

OpenStudy (anonymous):

Yes yes, I can always multiply either of them by -1 first :P :D

myininaya (myininaya):

but you didn't say that so you fail muhahaha

OpenStudy (anonymous):

lol, I left it for you to mention ...

myininaya (myininaya):

2x + 5y = -3 -15x - 5y = -10 ---------------- -13x+0=-13 -13x=-13 x=1

myininaya (myininaya):

Do you know how to find y?

myininaya (myininaya):

given x

OpenStudy (anonymous):

Now I know that you can add too :P

myininaya (myininaya):

I'm sometimes good at adding.